The width, w, of a rectangle is 4 cm shorter than the length,l. If the area of the rectangle is 60 centimeters squared, which equation models the situation correctly?
Answer by jorel555(1290)

If the length is l, then w, width, equals l-4. So your equation looks like:
A=l x w
60= l(l-4)
Solving, we get:
lē-4l-60=0
(l-10)(l+6)=0
l=10, -6
Your length is 10 cm, width is 6 cm. ☺☺☺☺